determine x-1/x+2 dx

Grade:

2 Answers

Shantnu Patlan
12 years ago

Here, you should subsitute (x+2).

Answer : (x+2)-3log(x+2)

Ashwin Muralidharan IIT Madras
12 years ago

Hi Seipati,

 

Here (x-1)/(x+2) = (x+2-3)/(x+2) = 1 - (3/[x+2])

 

So now integrate. That will be x - 3*ln(x+2) + Const.

 

Hope that helps.

 

Regards,

Ashwin (IIT Madras)